OpenStudy (anonymous):

Two urns each contain green balls and blue balls. Urn I contains 4 green balls and 6 blue balls, and Urn II contains 6 green balls and 2 blue balls. A ball is drawn at random from each urn. What is the probability that both balls are blue?

OpenStudy (anonymous):

The probability of drawing a blue ball from each urn is the probability of drawing a blue ball from the Urn 1 multiplied by the probability of drawing a blue ball from Urn 2.

OpenStudy (anonymous):

It is sort of like flipping 2 coins and asking for the chance of getting 2 heads. The chance of getting a head on Coin 1 is 1/2, and it's the same for Coin 2. The chance of getting heads on both coins is (1/2) * (1/2) = (1/4)

OpenStudy (anonymous):

But in this problem the probability is different for drawing a blue ball because there are different numbers of blue and green balls.

OpenStudy (anonymous):

A 2/51 B 3/20 C 1/10 D 4/153

OpenStudy (anonymous):

i still do not understand

OpenStudy (anonymous):

How many total balls in Urn 1? 6 blue + 4 green = 10 How many are blue? 6 blue So, the chance of drawing a blue ball out of Urn 1 is: (# of blue balls) / (total # of balls) = 6 / 10

OpenStudy (anonymous):

Does that help? Now you can do the same thing for Urn 2... find # of blue balls / total balls in Urn 2

OpenStudy (anonymous):

but the answer isn't up there ^

OpenStudy (anonymous):

what did you get for Urn 2?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

so do i multipy 6/10 and 1/4

OpenStudy (anonymous):

Sounds right to me. So the probability of 2 blues is (6/10) * (1/4) = (6 / 40)

OpenStudy (anonymous):

3/20

OpenStudy (anonymous):

right :) I didn't see it at first, but once you simplify the fraction, it's there :)

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

Glad to help :)